A block of mass 50kg is projected horizontally on arough horizontal floor .The coeffcient of friction is 0.1.The block strikes the light spring of stiffness k=100N/m with a velocity 2 m/s.The max compression of the spring is?
 
1)1  2)2  3)3  4)4 Please urgent want with solution

Grade:9

1 Answers

Vikas TU
7 years ago
From the energy conservation, (WET)
0.5kx^2 +  -0.1*500 = 0 – 0.5m*4
50x^2 = |50 – 100| 
50x^2 = 50
x = 1 meter